10. Find the equivalent resistance between points a and b for this resistor network. Draw the equivalent circuit and express the answer as Req = ‘ ar 30 Tao
Answer #1

Reg ((30 20160) 10) 80

here || means resistance in parallel and + means resistance in series

R_{eq} = (\frac{1}{\frac{1}{30}+\frac{1}{20}+\frac{1}{60}}+10)||80

R_{eq} = (\frac{1}{\frac{2+3+1}{60}}+10)||80

R_{eq} = (\frac{60}{6}+10)||80

R_{eq} = (10+10)||80

R_{eq} = \frac{20*80}{20+80}=16\Omega